OpenStudy (anonymous):

What is the solution of the following system of equations? y = (x+3)^2 - 4 y = 2x +5

OpenStudy (paxpolaris):

set them equal to each other and solve.

OpenStudy (paxpolaris):

\[\Large 2x+5=\left( x+3 \right)^2-4\]

OpenStudy (anonymous):

still don't get it.

OpenStudy (paxpolaris):

\[\large \implies 2x+5=\left( x^2+6x+9 \right) -4\]\[\large \implies 2x+5= x^2+6x+5\]

OpenStudy (paxpolaris):

move all terms to one side: \[\LARGE 0 = x^2+4x\] Solve for x's, and then get the corresponding y's

OpenStudy (anonymous):

OH. OKAY. THANK YOU SO MUCH :) YOU JUST SAVED ME <3

OpenStudy (paxpolaris):

\(\Large(x,y)= (0,5)\) or \(\Large(x,y)=(-4, -3)\)

OpenStudy (anonymous):

YUP, THANK YOU :)